### v1

@@ Abstract
This paper presents a method for accurately predicting the full distribution of stock returns, given a comprehensive set of 194 stock characteristics and market variables. Such distributions, learned from rich data using a machine learning algorithm, are not constrained by restrictive model assumptions and allow the exploration of non-Gaussian, heavy-tailed data and their non-linear interactions. The method uses a two-stage quantile neural network combined with spline interpolation. The results show that the proposed approach outperforms alternative models in terms of out-of-sample losses. Furthermore, we show that the moments derived from such distributions can be useful as alternative empirical estimates in many cases, including mean estimation and forecasting. Finally, we examine the relationship between cross-sectional returns and several distributional characteristics. The results are robust to a wide range of US and international data.
Keywords: Distribution forecasting, Cross-section of stock returns, Anomalies, Quantile forecasting, Machine learning, Neural networks, Splines, International markets
@@ 1 Introduction
Despite considerable scepticism that returns are inherently difficult to predict (Stock and Watson, 2017), a substantial body of evidence has emerged over several decades showing that the first two moments of stock return distributions can be predicted to some extent using economic variables (Ang et al., 2006; Ang and Bekaert, 2007; McLean and Pontiff, 2016). As a practical matter, mean-variance analysis provides investors with a framework for making informed choices if they accept the constraints imposed by assumptions such as multivariate normality of stock returns or a quadratic utility function. At the same time, economic agents engaged in a wide range of activities, including risk management, portfolio selection, derivatives pricing and asset pricing, have a much more complex set of preferences. They are endowed with asymmetric utilities such as disappointment aversion (Gul, 1991), general preferences (Kimball, 1993), quantile preferences (Manski, 1988) or prospect theory (Kahneman and Tversky, 1979), which require distributional information beyond the first two moments. A complete distribution forecast is necessary to fully characterise the uncertain future evolution of returns for such decision-makers. A major challenge is to develop effective forecasting techniques for a return distribution, especially given the limitations of available data.
Forecasting stock return distributions is challenging due to a number of factors, including low signal-to-noise ratio, the presence of fat tails, volatility clustering and the potentially high-dimensional nature of the relevant information sets. Worryingly, much of the existing work relies on restrictive models and assumptions, does not take into account the possible non-linear interaction of a large number of variables, and does not take into account well-known features of the data, such as non-Gaussianity and heavy tails. With rapid improvements in the accessibility and availability of large datasets and computer science algorithms, it is tempting to ask whether the problem can be significantly improved using methods that focus on learning patterns from data. In line with recent efforts to move away from exclusive reliance on models to machine learning approaches (Athey and Imbens, 2019; Mullainathan and Spiess, 2017), we propose to use machine learning to predict the full distribution of out-of-sample stock returns. We provide ample evidence that such an approach makes sense using a wide range of US and international stocks.
Our contributions can be divided into two distinct areas. First, we present a novel distribution forecasting method based on the combination of a two-stage quantile neural network with cubic B-splines interpolation. The two-stage quantile neural network uses a substantial number of stock characteristics and market variables in two sub-networks to predict a set of $\tau$-quantiles of returns, which are then interpolated to obtain the full distribution. This crucial step allows the derivation and use of key moments of the distribution, including mean, volatility, skewness and kurtosis. The model has been trained on US data to demonstrate out-of-sample performance from both a time perspective, using a rolling window in the US, and a cross-sectional perspective on international stocks. The performance of the proposed method is compared with that of traditional parametric models and other neural network models. The results show that the proposed approach outperforms the benchmark models in terms of both out-of-sample average quantile loss and higher out-of-sample $R^{2}$ for mean forecasts, as well as lower out-of-sample error in variance forecasts. It is crucial to highlight that utilising information from the entire distribution enhances the forecasting accuracy of the mean and variance.
Second, the paper examines how different distributional characteristics are reflected in the cross-section of stock returns. A comprehensive dataset covering both US and international stocks is used in this study, with results presented for individual regions and for the full and liquid samples. In particular, we examine the influence of individual $\tau$-quantiles as well as mean, volatility, skewness and kurtosis derived from the predicted distribution. The profitability of long-short portfolios constructed on the basis of $\tau$-quantile forecasts of individual stocks is analysed. It is shown that central $\tau$-quantiles are priced into the cross-section of stock returns. There is a significant decrease in the profitability of long-short portfolios based on $\tau$-quantiles as we move from the centre of the distribution towards the tails. We show that the mean obtained by numerical integration of the interpolated distribution forecast leads to significantly higher out-of-sample long-short portfolio returns than the mean directly predicted by the competing models. In addition, we show that no other moment studied is significantly monotonically priced in the cross-section of stock returns. This provides additional support for previously conflicting results on the role of variance, skewness and kurtosis in asset pricing.
In order to illustrate the efficacy of our methodology, we present the forecasted probability function obtained using our methods on the case of the Apple stock. Figure 1 illustrates the evolution of the forecasted distribution of the 22-day-ahead returns over the period 1996-2020. The distribution is obtained without any model assumptions, utilising a substantial number of characteristics. Yet, it captures the dynamics precisely.
Figure 1: Example of forecasted probability density function over time. This figure displays the forecasted probability density function for out-of-sample 22 days-ahead returns of an Apple stock (CUSIP 03783310) generated using our two-stage quantile neural network model.
In our methodology, we prioritise the forecasting of quantiles and subsequently approximate the distribution function rather than directly estimating a parametric distribution. In light of the stylised facts of stock returns, including the presence of fat tails, asymmetry and a domain extending from $-1$ to +$\infty$+ ∞, it is challenging to specify a distributional form that accurately captures these characteristics. For example, the normal distribution is unable to capture the fat tails, the Student-t distribution is unable to capture the asymmetry, and the skewed Student-t distribution is unable to capture the domain restriction. Conversely, our approach approximates the distribution function in a data-driven manner, eschewing the imposition of strong parametric assumptions. Furthermore, our method is capable of estimating discrete probabilities, such as the probability mass at -1, which is of particular importance for long-horizon return forecasts. The selected methodology yields intriguing outcomes, such as the predicted discrete probability mass at 00 for certain stocks, which would be unattainable through the parametric distribution function.
The empirical results presented in this study are based on a comprehensive analysis of both U.S. and international stock data spanning the period from 1995 to 2018. To ensure the robustness of our results, we focus on the full sample as well as a restricted liquid sample. To demonstrate the generalisability of our distribution forecasting method, we train the model using only U.S. data and then apply it to international stocks and U.S. stocks in an out-of-sample manner. All requisite stock-level characteristics and market variable definitions are consistent across the United States, thereby eliminating the potential for country-specific variables (e.g., macroeconomic variables) or variables available only for a subset of stocks (e.g., options-based variables) to influence the model. This consistency allows for the transfer of predictability to the international domain.
By focusing on the aspect of predictability, we demonstrate that our distribution forecasting method exhibits a notable degree of outperformance in terms of average out-of-sample quantile loss when compared to forecasts generated by the traditional time series models as well as complex neural networks. Furthermore, the distribution forecast provides substantial benefits when forecasting the mean of stock returns. The mean forecast obtained by means of numerical integration of the interpolated distribution forecast gives rise to significantly higher out-of-sample $R^{2}$ values than the mean forecast derived directly from neural network models using mean squared error loss. With regard to volatility forecasting, the volatility forecast derived from numerical integration of the interpolated distribution forecast results in a reduction in the out-of-sample mean absolute deviation and root mean squared error in comparison to those forecasted from the GARCH model. These findings are consistent across all regions and both full and liquid samples.
In light of the distribution forecasting capability of our model, we undertake a further empirical investigation into the role of different distributional characteristics in an asset pricing application. This study examines the profitability of long-short portfolios formed based on $\tau$-quantile forecasts of individual stocks, as well as the roles of mean, volatility, skewness, and kurtosis. Our findings indicate that, among all the quantiles, only the central quantiles are reflected in the cross-section of stock returns. Furthermore, we observe a notable decline in the profitability of long-short portfolios formed on these quantiles as we move away from the central region of the distribution towards the tails. Furthermore, we examine the profitability of long-short portfolios based on the mean versus the median forecast. Our findings indicate that the mean forecast leads to significantly higher out-of-sample long-short portfolio returns on the full sample. However, this is not the case on the liquid sample, which highlights the importance of underlying distributional asymmetry. Finally, we demonstrate that no other studied moment is priced in the cross-section of stock returns, providing further evidence to resolve the previously observed conflicting results on the roles of variance, skewness, and kurtosis in asset pricing. The following section will present the results of the analysis.
1.1 Relation of our work to the literature
Our work relates to several strands of the literature. We contribute to stock return forecasting, which is a central topic in the asset pricing literature with a long history. In recent years, machine learning methods have become a tool of choice in this literature due to their ability to capture complex non-linear relationships, handle high-dimensional data with low signal-to-noise ratios, and provide superior sample-to-sample performance compared to traditional approaches and superior out-of-sample performance compared to more traditional approaches (Gu et al., 2020). For example, Bryzgalova et al. (2020) shows how to build better test assets by creating managed portfolios using decision trees, Dong et al. (2022) links the cross-sectional predictability of stock returns with the time-series predictability of aggregate market returns by using anomaly portfolios to predict market returns, Gu et al. (2021) uses autoencoders to build a factor asset pricing model that allows for non-linear factor loadings. Successful applications of machine learning techniques have also been demonstrated in other domains and asset classes, such as the prediction of bond returns by Bianchi et al. (2021), the ex-ante identification of good and bad performing mutual funds by Kaniel et al. (2023), the prediction of option returns by Bali et al. (2023), or the use of convolutional neural networks to use price images to predict stock returns by Jiang et al. (2023). Furthermore, the robustness of many machine learning applications in stock return forecasting has been confirmed in an international context by Azevedo et al. (2023), Cakici et al. (2023) and Tobek and Hronec (2021).
The literature on forecasting the distribution of stock returns is more limited. Our work is most closely related to Liu (2023) in terms of forecasting the quantiles of future stock returns in the cross-section. Liu (2023) follows Gu et al. (2020) and compares a variety of linear models and machine learning methods in a quantile regression setting. [^1] Models compared include linear quantile regression, elastic net quantile regression, quantile random forests, quantile gradient boosted regression trees, and feed-forward neural networks. In terms of out-of-sample prediction performance, he shows that quantile neural networks significantly outperform other models. He introduces a quantile-based risk premium measure as a weighted average of five predicted quantiles, where the weights correspond to an area under an empirical step function. He also constructs a robust measure of volatility (interquantile range) and skewness from the $25\%$, $50\%$ and $70\%$ quantile forecasts, documenting that skewness is priced into the cross-section of stock returns. Our paper differs from Liu (2023) in several ways.
We present a novel neural network architecture for quantile forecasting and show how a list of quantile forecasts can be interpolated into a dense representation of the distribution function. Given our denser representation of the distribution, we derive moments using numerical integration, which provides a more accurate representation of the moments than directly using robust quantile-based measures. We also show how to adjust the predicted higher moments to account for the lack of extreme observations in the tails of the distribution. In addition, we provide international evidence for models trained on US data only and show that our approach generalises well to international stocks. Our results are consistent with the importance of obtaining the mean prediction from the distribution [^2] Even if the implementation differs. rather than directly using the mean squared error; however, we find no evidence that skewness is priced into the cross-section of stock returns, either in the US or internationally. We also examine the role of individual $\tau$-quantiles in the cross-section of stock returns.
Another related paper is Yang et al. (2024), which builds on Gu et al. (2021) by using a conditional autoencoder model to approximate the distribution of future stock returns via a step function using a conditional quantile variational autoencoder (CQVAE). The authors show that the CQVAE model effectively captures the non-linear dependencies between stock returns and latent factors, leading to superior out-of-sample forecasting performance compared to traditional linear models. They also show that the mean returns estimated from the step-function approximation of the distribution function are more accurate and lead to better portfolio performance. While both our paper and Yang et al. (2024) use machine learning to predict the distribution of stock returns, there are significant differences. The CQVAE model uses latent factors learned by a variational autoencoder, while our approach uses stock characteristics and market variables directly as predictors. We go further by deriving key distributional moments such as mean, volatility, skewness and kurtosis through numerical integration, in contrast to the CQVAE model’s focus on using quantile forecasts primarily for mean estimation. In addition, we validate the generalisability and robustness of our model using both US and international stock data, demonstrating its broad applicability, whereas Yang et al. (2024) focuses solely on US stocks.
The distributions of individual stock returns are relevant to the asset pricing literature, as can be seen from the from the number of anomalies based on different distribution characteristics, see Hou et al. (2020). However, these anomalies are typically based on historical sample estimates rather than out-of-sample forecasts, leading to significant estimation errors, especially when the persistence of tail-based distributional of tail-based distributional properties is low. For example, estimating the third and fourth moments rather than the first two moments Kim and White (2004). A promising alternative to distributional forecasting is the use of options to derive implied moments (An et al. (2014), Huang and Li (2019), Alexiou and Rompolis (2022)), although this method is limited by the coverage and liquidity of the and liquidity of options markets, especially outside the US. Instead of relying on historical sample estimates or option data, we use distributional characteristics derived from our out-of-sample distribution forecasts.
A large body of literature examines how distributional characteristics affect the cross-sectional distribution of stock returns. of stock returns. There is a consensus on the ambiguous relationship between realised volatility and future stock returns. and future stock returns (Bollerslev et al. (2020), Ang et al. (2006)). We contribute further evidence on the lack of a clear relationship between forecast volatility and future stock returns. stock returns. A similar lack of documented relationships also exists for kurtosis, although the the literature here is less extensive (Bollerslev et al., 2020).
The role of skewness, on the other hand, presents a more complex picture. Skewness is crucial in understanding individual stock returns, as Bessembinder (2018) notes that the positive skewness of individual stock returns fundamentally explains why a small fraction of stocks generate long-term stock market wealth. A series of papers, extensively reviewed by Bali et al. (2016), examine the risk premium associated with the cross-sectional skewness of stock returns. of stock returns.
In addition to asset pricing, the importance of predicting distributions is evident in risk management, where the left tail of the distribution is of particular interest. Value at Risk (VaR) is a standard risk management tool and corresponds to the $\tau$ quantile of the conditional distribution of an asset’s return. Various approaches to VaR estimation, ranging from non-parametric (historical simulation) to semi-parametric to parametric, have a long history in the literature. Unlike non-parametric estimation, parametric VaR estimation requires assumptions about specific return distributions and model dynamics. The foundation of this approach is based on the GARCH volatility model of Engle (1982) and Bollerslev (1986), which is typically used together with the assumption that the residuals are distributed according to a specific distribution, typically normal or Student-t. Among semi-parametric estimation approaches, a prominent example is the CaViaR model of Engle and Manganelli (2004), which uses quantile regression to directly estimate the model dynamics without making further assumptions about the error distribution. See Nieto and Ruiz (2016) for a comprehensive review of VaR estimation methods.
While our work and the VaR estimation literature are clearly related in that both rely on accurate quantile forecasts, there are also obvious differences. First, we focus on the entire cross-section of stock returns, while the VaR estimation literature tends to work with a single asset, usually representing the entire portfolio. This means that we can rely on a rich set of information in the form of stock-level characteristics, whereas the VaR estimation literature must rely mainly on the historical returns of the asset itself. Second, our ultimate goal is to forecast the full distribution of stock returns, not just specific $5\%$ or $1\%$ quantiles. As a result, we can derive any distributional risk measure, such as conditional value at risk (CVaR), skewness or kurtosis.
@@ 1.1 Relation of our work to the literature
The rest of the paper is organised as follows. Section 2 introduces our distribution forecasting method, including quantile neural networks and subsequent interpolation of the quantiles into a full distribution function using splines. Section 3 presents the empirical results, consisting of an out-of-sample assessment of predictability and an analysis of the asset pricing implications of various distributional features.
@@ 2 Two-stage quantile neural networks
Let the returns $r_{i,t}$ collected over $t=1,\dots,T$ months and $i=1,\dots,N_{t}$ individual stocks be random variables with a distribution function [^3] In the case of the forecast distribution of stock returns, $Q_{r_{i,t}}(0)=-1$ and $Q_{r_{i,t}}(1)=\infty$. $F(r_{i,t})$. We are interested in forecasting a set of multiple $\tau$ quantiles $Q_{r_{i,t+1}}^{\mathcal{T}}=\Big{\{}Q_{r_{i,t+1}}(\tau_{1}|\mathcal{F}_{t}),\ldots,Q_{r_{i,t+1}}(\tau_{K}|\mathcal{F}_{t})\Big{\}}$ for $\forall\tau\in\mathcal{T}=\{\tau_{1},\ldots,\tau_{K}\}$ where $Q_{r_{i,t}}(\tau\big{|}\mathcal{F}_{t})=\inf\{r_{i,t}\in\mathbb{R}:F(r_{i,t})\geq\tau|\mathcal{F}_{t}\}$ of a stock return at time $t+1$ is conditional on the information set $\mathcal{F}_{t}$ available at time $t$.
Forecasting multiple $\tau$-quantiles is possible with a multi-output feed-forward neural network model based on aggregating individual $\tau$-quantile losses into a single loss function (as described later in this section). In addition to being computationally more efficient than training individual models for each $\tau$, joint training for all $\tau$ quantiles allows the model to capture the relationship across the entire distribution. As a side effect, even without introducing penalisation for quantile crossing [^4] Quantile crossing refers to the situation where $Q(\tau_{l})>Q(\tau_{h})$ when $\tau_{l}<\tau_{h}$, violating the quantile monotonicity condition., the model naturally captures the monotonicity requirement of quantile predictions since there are no quantiles that are crossed in our quantile forecasts.

### v2

@@ Abstract
We propose a novel machine learning approach for forecasting the distribution of stock returns using a rich set of firm-level and market predictors. Our method combines a two-stage quantile neural network with spline interpolation to construct smooth, flexible cumulative distribution functions without relying on restrictive parametric assumptions. This allows accurate modelling of non-Gaussian features such as fat tails and asymmetries. Furthermore, we show how to derive other statistics from the forecasted return distribution such as mean, variance, skewness, and kurtosis. The derived mean and variance forecasts offer significantly improved out-of-sample performance compared to standard models. We demonstrate the robustness of the method in US and international markets.
Keywords: Distribution forecasting, Cross-section of stock returns, Anomalies, Quantile forecasting, Machine learning, Neural networks, Quantile neural networks, Splines, International markets, Skewness, Kurtosis
@@ 1 Introduction
Predicting the distribution of stock returns has become increasingly important because modern financial decision-making requires more than just point estimates. Traditional approaches that focus on predicting the mean or variance of returns impose restrictive assumptions, such as multivariate normality or quadratic utility, that rarely hold in practice. These methods fail to capture the complexity of investor preferences, which depend on higher order moments and distributional properties such as asymmetry, tail behaviour and discontinuities (Gul, 1991; Kimball, 1993; Manski, 1988; Kahneman and Tversky, 1979), as well as features of the data such as non-Gaussianity and heavy tails. To address these limitations, we develop a machine learning-based framework for forecasting the distribution of stock returns using a rich set of firm-level and market variables.
Specifically, this paper presents a novel approach that uses a two-stage quantile neural network combined with spline interpolation to generate predictive distributions. Our method flexibly models non-Gaussian, heavy-tailed return distributions without relying on parametric assumptions. It captures non-linear interactions among a large number of predictors, allows estimation of key distributional moments, and provides accurate forecasts in high-dimensional settings with low signal-to-noise ratios. Our approach performs well in both US and international markets and delivers significant improvements in out-of-sample loss metrics compared to existing parametric and machine learning benchmarks in the prediction of next month returns.
The ability to forecast next month returns distributions rather than isolated moments has a number of key benefits. It supports decision frameworks that require knowledge of tail risk or asymmetry - such as those used in portfolio optimisation, risk management and asset pricing. It improves the estimation of standard moments (mean, variance, skewness, kurtosis) through numerical integration, which we show improves predictive performance over direct moment forecasting. Finally, it also enables new empirical investigations into the pricing of distributional characteristics across assets.
We contribute to the literature in two main ways. First, we propose a new method for distribution forecasting that combines quantile neural networks with cubic B-spline interpolation to construct smooth and flexible cumulative distribution functions. Unlike existing approaches, our method is non-parametric, data-driven and generalises well across the globe.
Second, we show how distributional forecasts can be used to improve asset pricing applications. Specifically, we examine the pricing of predicted quantiles and moments in the cross-section of stock returns. We show that the mean return derived from the interpolated distribution significantly outperforms direct mean forecasts in out-of-sample portfolio returns. Moreover, we find that only the central quantiles are systematically priced, with the profitability of long-short portfolios declining as one moves towards the tails. Contrary to some previous findings, we find no consistent pricing of volatility, skewness or kurtosis using our distribution-derived estimates. These results provide new insights into the role of higher-order moments in asset pricing and help reconcile previous mixed evidence in the literature (Bollerslev et al., 2020; Ang et al., 2006; Bali et al., 2016).
Our work builds on and extends a growing body of research applying machine learning to asset pricing and return forecasting. Recent studies have demonstrated the superiority of machine learning methods in capturing complex non-linear patterns and improving out-of-sample prediction accuracy (e.g. Gu et al. (2020a, 2021); Bryzgalova et al. (2020); Tobek and Hronec (2021); Dong et al. (2022)). However, most of these studies focus on predicting mean returns or linear factor models. In contrast, our approach targets the entire distribution and goes beyond previous quantile prediction methods by interpolating across a dense grid of quantiles to construct smooth distributions.
We also complement recent contributions to distributional forecasting in asset pricing. For example, Liu (2023) demonstrates the usefulness of quantile regression for forecasting returns and constructing quantile-based risk premia. Our approach differs in that we interpolate a richer set of quantile forecasts onto a denser representation of the distribution and derive distributional moments via numerical integration rather than relying on pre-specified interquantile ranges. We also show how to adjust the predicted higher moments to account for the lack of extreme observations in the tails of the distribution. Furthermore, our method generalises to international markets and is validated using models trained only on US data.
Another related strand is the work of Yang et al. (2024), who use a conditional quantile variational autoencoder (CQVAE) to approximate future return distributions. While both approaches use machine learning to generate distribution forecasts, our method uses observable firm-level predictors and explicit quantile targets, in contrast to the latent factor representations used in CQVAE. In addition, we provide a detailed decomposition of distributional moments and their pricing implications for both US and global equity markets. In contrast, Yang et al. (2024) constructs quantile forecasts with a primary focus on mean estimation.
Our contribution is also relevant to the literature on distributional properties in asset pricing (Hou et al., 2020; Bessembinder, 2018; Kim and White, 2004; An et al., 2014; Huang and Li, 2019; Alexiou and Rompolis, 2022). Existing evidence on the pricing of volatility, skewness and kurtosis is mixed and often relies on ex-post sample estimates or option-implied measures, which have limited availability. Our model provides an alternative by using out-of-sample distribution forecasts based on observable predictors to study the pricing of distributional characteristics. We show that only the mean and certain central quantiles carry significant pricing information, while higher moments do not appear to be systematically priced.
In summary, this paper presents a novel machine learning-based method for forecasting stock return distributions and provides empirical evidence on its predictive accuracy and relevance for asset pricing. Our results underscore the value of distribution forecasting in financial applications and suggest that incorporating richer distributional information can improve return forecasting and portfolio performance.
@@
The rest of the paper is organised as follows: Section 2 presents the distributional forecasting methodology, including quantile neural networks and spline interpolation. Section 3 describes the empirical setup and data, reports results on out-of-sample forecasting performance and asset pricing applications. Section 4 concludes.
@@ 2 Two-stage quantile neural networks
Let the returns $r_{i,t}$ collected over $t=1,\dots,T$ months and $i=1,\dots,N_{t}$ individual stocks be random variables with a cumulative distribution function (CDF) [^1] In the case of the forecast distribution of stock returns, $Q_{r_{i,t}}(0)=-1$ and $Q_{r_{i,t}}(1)=\infty$. $F(r_{i,t})$. Let $\tau_{1},\dots,\tau_{K}\in(0,1)$ denote the quantile levels of interest, and define the set $\mathcal{T}=\{\tau_{1},\ldots,\tau_{K}\}$. We are interested in forecasting the corresponding conditional quantiles $Q_{r_{i,t+1}}^{\mathcal{T}}=\big{\{}Q_{r_{i,t+1}}(\tau_{1}|\mathcal{F}_{t}),\ldots,\\ Q_{r_{i,t+1}}(\tau_{K}|\mathcal{F}_{t})\big{\}}$, where each quantile is defined as
$Q_{r_{i,t}}(\tau\mid\mathcal{F}_{t})=\inf\left\{r\in\mathbb{R}:F(r)\geq\tau\mid\mathcal{F}_{t}\right\},$
that is, the $\tau$-quantile of the return distribution at time $t+1$ conditional on the information set $\mathcal{F}_{t}$ available at time $t$. Forecasting multiple $\tau$-quantiles, where $\tau\in(0,1)$ is possible with a multi-output feed-forward neural network model based on aggregating individual $\tau$-quantile losses into a single loss function (as described later in this section). In addition to being computationally more efficient than training individual models for each $\tau$, joint training for all $\tau$ quantiles allows the model to capture the relationship across the entire distribution. As a side effect, even without introducing penalisation for quantile crossing [^2] Quantile crossing refers to the situation where $Q(\tau_{l})>Q(\tau_{h})$ when $\tau_{l}<\tau_{h}$, violating the quantile monotonicity condition., the model naturally captures the monotonicity requirement of quantile predictions since there are no quantiles that are crossed in our quantile forecasts.